# Visibly Fair Mechanisms

### v1

@@ Abstract
Priority-based allocation of individuals to positions are pervasive, and elimination of justified envy is often, an absolute requirement. This leaves serial dictatorship (SD) as the only rule that avoids justified envy under standard direct mechanisms. What if SD outcomes are undesirable from a designer’s perspective? We propose visible fairness, which demands fairness relative to the (potentially purposefully incomplete) preference information the mechanism elicits. Visibly fair mechanisms generalize SD; we fully characterize them and provide necessary and sufficient conditions for strategy-proofness. We show how to apply these results to design strategy-proof visibly fair rules that satisfy a broad class of distributional objectives. Visible fairness, however, results in a new information-efficiency trade-off: meeting distributional goals leads to the avoidance of eliciation of information about preferences that could prevent inefficiencies.
@@ 1 Introduction
Priority-based assignments are pervasive in a wide range of real-world matching contexts, including university admissions, public-sector placements, and cadet-branch allocations in military academies (see, e.g., Balinski and Sönmez, 1999; Sönmez and Switzer, 2013). In a typical priority-based system, participants receive a strict ranking—such as exam scores or a merit list—and are assigned to positions accordingly. The central fairness requirement in these contexts is that no lower-ranked participant should occupy a seat that a higher-ranked participant strictly prefers; otherwise, the latter has a legitimate grievance, known as justified envy (Abdulkadiroğlu and Sönmez, 2003). Such fairness concerns become even more pronounced when the priorities at stake represent strongly protected interests—like property rights or national exam rankings—where even a single instance of justified envy can trigger legal and administrative challenges. [^1] One example of these legal challenges is the Federal University of Bahia hiring suit (Brazil, 2025), where a federal judge blocked the university from hiring a lower-scoring quota applicant and ordered the single vacancy awarded to the exam’s top scorer (Agência Estado, 2025). Another was the Italian national residency “fiasco”, in which the Regional Administrative Tribunal of Lazio annulled a ruling that forced higher-ranked doctors to forfeit more-preferred specialties while lower-ranked peers advanced (Focus.it, 2016). Both rulings treat the harm as a breach of the merit order—i.e., a violation of justified envy—showing that such breaches readily provoke litigation.
Under the standard approach of designing direct mechanisms—where each participant reports a complete ranking over positions—Serial Dictatorship (SD) is in fact the only mechanism that can satisfy non-wastefulness and no-justified-envy. In SD, the highest-priority agent picks their top choice, the next agent picks from the remaining positions, and so on. This procedure prevents any lower-ranked agent from ending up in a spot that a higher-ranked agent strictly prefers, ensuring no justified envy. However, SD can produce allocations that are misaligned with policy goals, such as excessive clustering of top-ranked agents in a small set of elite locations, or an undesirable regional or demographic distribution of assignees. When considering standard direct mechanisms, there is no alternative to SD that both respects strict priorities and operates on full preference lists. Thus, policymakers appear to face a dilemma: given the requirement for “fairness by priority”, how can one construct rules in pursuit of better distributional outcomes?
@@
1. In the Indian Administrative Services (IAS), officers were formerly assigned to state cadres in a priority-driven process aligned with exam-based merit. This arrangement, which was essentially a serial dictatorship with some modifications, produced undesirable allocations exhibiting homophily, i.e., a propensity for officers to serve in or near their home regions. Such geographic clustering compromised the national integration objective of the service (Thakur, 2023). A 2017 reform imposed a zone-based scheme: all cadres were partitioned into five geographic zones, and each officer now submits a separate ranking of cadres within every zone rather than a single list over the entire country (see Figure1). The revised mechanism guarantees that no lower-ranked officer receives a cadre preferred by a higher-ranked officer inside the zone that ultimately matches them, while the zonal structure itself disperses high-ranked officers more evenly across India—advancing diversity goals without overriding the preferences participants actually report.
@@
$>$Zone 2 Uttar Pradesh
@@
$>$Zone 5 Andhra Pradesh
@@
2.  Under the U.S. Military Academy matching process to military branches used in 2006, each cadet (i) ranked the branches and (ii) stated, for every branch, whether they would accept a longer service obligation in exchange for a priority boost (Sönmez and Switzer, 2013). This elicitation did not allow them to report cross-branch trade-offs. Omitting this information let the Academy honour the official order-of-merit list while making it possible to steer more cadets into longer commitments that would have been rejected under full preference elicitation. As Sönmez (2024) explains, this design kept such priority violations hidden:
@@
As the preceding cases show, limiting what participants may report can hide genuine priority breaches. We say a mechanism is visibly fair when, given the elicited (partial) preferences, no outcome appears to violate priority. For example, if seats are partitioned into zones and agents may only rank seats within each zone, allocation by global priority inside every zone looks perfectly fair—even though a cross-zone comparison (never elicited) might reveal a lower-ranked agent holding a seat a higher-ranked agent prefers. By restricting the scope of reported preferences, designers can pursue goals such as geographic diversity while keeping any latent violations invisible.
For example, if positions are partitioned into zones and agents may only rank positions within each zone, an allocation by priority inside every zone looks perfectly fair—even though a cross-zone comparison (never elicited) might reveal a lower-ranked agent holding a seat a higher-ranked agent prefers. By restricting the scope of reported preferences, designers can pursue goals such as geographic diversity while keeping any underlying violations of priorities invisible.
Inspired by these observations, we examine the design problem of assigning participants to seats under a strict priority ordering while maintaining visible fairness. In contrast to standard models that fix a preference-reporting format, in this framework the designer chooses both the message space (the form of partial preferences agents can report) and the outcome rule. Our analysis provides a framework and results on how to configure these two elements together so as to achieve desired policy objectives.
@@
Our analysis delivers three main sets of results. First, we pin down the precise structure that visible fairness imposes on allocation rules. Theorem 1 shows that any visibly–fair mechanism must operate as an $m$-queue allocation: officers are processed in strict priority order, and each is assigned a state that is undominated within the partial ranking she is allowed to report. When the message space is zonal, this characterization specializes in Theorems 2 and 3: the only visibly–fair rules are partitioned (and, with cross-zone rankings, ranked–partitioned) priority mechanisms. While visible fairness implies serial dictatorship (and therefore strategy-proofness) when using direct mechanisms (Corollary 2), that is not the case for general message spaces. We we show in Theorem 4 that strategy-proofness is obtained exactly when the mechanism also satisfies expressiveness and (weak) availability, two properties that rule out profitable deviations when the mapping from message profiles to outcomes are more general.
Second, we introduce a flexible way to encode distributional goals through modular upper-bounds. A quota system is modular when every bound caps a single type within an arbitrary subset of states (Definition 10). Modular bounds induce a natural partition of states—formally captured in Definition 13—that yields a message space where all states subject to the same collection of caps fall in the same “zone.” Building on this structure, the Modular Priority Mechanism (Definition 14) processes officers by priority while dynamically blocking zones whose relevant caps have just filled. Theorem 6 proves that this mechanism simultaneously respects every modular bound, remains visibly fair, and is strategy-proof.
Yet, as Example 6.1 and the impossibility result in Theorem 7 show, no static (one-shot) rule can simultaneously (i) satisfy visible fairness, (ii) respect modular upper-bounds, and (iii) not be Pareto-improved by quota-feasible allocations. The root of the conflict is informational. To guarantee that every admissible report keeps the quotas intact, the designer must prune the message space in advance, excluding comparisons whose truthful revelation could otherwise force a violation. This secures fairness and feasibility but withholds preference information that would uncover (and execute) mutually beneficial swaps still compatible with the same caps.
@@
A foundational theory for mechanisms with restricted message spaces comes from Green and Laffont (1986), who study settings where agents are restricted to a limited message space that depends on their true state. Their key insight is that by constraining the information an agent can reveal, the set of implementable outcomes can be expanded beyond what is possible in standard direct mechanisms. We deliberately design message spaces to elicit only partial preferences and make certain comparisons impossible, enabling the achievement of distributional goals that would otherwise conflict with fairness claims.
@@
Beyond the already discussed cases of the Indian Administrative Service (IAS) cadre allocation (Thakur, 2021), the U.S. Military Academy’s cadet-branch matching (Sönmez and Switzer, 2013; Sönmez, 2013), and Chinese college admissions (Hu et al., 2025), many real-world mechanisms also limit the extent of preference reporting. In school choice, for example, some systems cap the number of schools an applicant may rank (Haeringer and Klijn, 2009a; Calsamiglia et al., 2010), while others allow students to “bundle” schools into groups without inter-group comparisons (Huang and Zhang, 2025).
@@
Dynamic procedures also restrict preferences and by doing so, might reduce complexity (Pycia and Troyan, 2023). Bó and Hakimov (2022) propose Iterative Deferred Acceptance by letting agents choose from menus, obtaining stable results without demanding complete rankings, and Bó and Hakimov (2024); Mackenzie and Zhou (2022) extend this idea to mechanisms that sequentially offer feasible outcomes—enhancing privacy and performing well in controlled experiments. Even small constraints, like limiting the length of rank-ordered lists, can disrupt classical incentive properties: Haeringer and Klijn (2009b) show how capping the number of ranked schools compromises the usual strategy-proofness of the Deferred Acceptance procedure. Meanwhile, Caspari and Khanna (2025) propose precise conditions for stability and incentives with non-standard preference formats. Collectively, these studies highlight how restricting preference articulation can open new design possibilities while preserving key desiderata—an insight we leverage in defining and deploying visible fairness.
Decerf et al. (2024) is perhaps the most conceptually related work to ours. Their notion of incontestable assignments describes an environment where students cannot fully observe others’ preferences or placements, leaving them unable to identify certain envy issues. This parallels our concept of visible fairness, in which certain violations become undetectable. The key difference, however, is that Decerf et al. (2024) derive their informational constraints from the students’ limited ability to view the full outcome, whereas in our framework, these constraints are intentionally designed by the mechanism planner. Specifically, in our setup the designer restricts what agents can report so as to preclude distributional tensions that might otherwise manifest as visible grievances. In addition, while Decerf et al. (2024) accommodate a variety of school-specific priorities, our model considers a single, strict priority ranking that orders all participants.
@@
Partial preferences has been used in other contexts as well in market design. One strand lets participants declare indifference classes directly: Erdil and Ergin (2017); Manjunath and Westkamp (2021); Andersson et al. (2021) build mechanisms that treat weak orders—strict ranks punctuated by ties—as the primitives, and then exploit those ties to recover efficiency and strategy‐proofness. A second strand considers problems in which not every pair of outcomes can be compared, evaluating which adaptation of standard properties, such as stability, can nonetheless sustain strategy-proofness under suitable conditions (Caspari and Khanna, 2025; Kuvalekar, 2023). Typically, these frameworks rely on “weak stability,” where agents who are indifferent or indecisive simply cannot block assignments. In contrast, our approach presumes agents do have complete preferences but are deliberately constrained from revealing them in full.
@@
At the core of this field lies the tension between rigid distributional quotas and the flexible preferences of market participants. Fragiadakis and Troyan (2017) demonstrated this through military cadet matching, where static reservation systems created inefficiencies by locking seats for specific groups prematurely. Their dynamic quota mechanism represented a paradigm shift, adjusting reservation targets based on revealed preferences while maintaining strategy-proofness. This approach inspired subsequent innovations like the Adaptive Deferred Acceptance (ADA) mechanism by Goto et al. (2017), which introduced hereditary constraints—rules where satisfying a constraint automatically satisfies all its subsets. The ADA mechanism’s success in Japanese medical residency matching showed that carefully designed constraints need not sacrifice core market principles like nonwastefulness and strategy-proofness.
We add to this literature by demonstrating how appropriately crafted message spaces—which limit how agents can express their preferences—allow mechanism designers to achieve distributional objectives while preserving fairness and incentive properties.
A significant direction for future research is to design message spaces that can accommodate floor constraints, where balancing stability and efficiency remains a central theoretical challenge. Tomoeda (2018) established mathematical conditions where type-specific minimum quotas could coexist with stable matchings, requiring that institutions’ preferences satisfy substitutability across types—a condition frequently violated in practice. Ehlers et al. (2014) proposed softening this tension through probabilistic constraints, where schools could satisfy diversity targets over multiple matching cycles rather than single instances. This “soft bounds” approach proved particularly valuable in school choice contexts, where rigid racial or socioeconomic quotas often face legal challenges. Akin (2021) introduced a stability concept accommodating strict floor constraints, along with a mechanism that is strategy-proof for doctors and results in a stable matching.
Structure of the paper. In Section 2, we introduce the model and definitions, covering partial preferences, feasible allocations, and our message-space framework. Section 3 then characterizes visibly fair mechanisms, identifying them as queue-allocation variants and examining applications such as zonal message spaces. Section 4 turns to incentives, specifying exactly when these mechanisms are strategy-proof via conditions of expressiveness and availability. Sections 5 and 6 present our central results on distributional objectives, including modular upper-bounds and the Modular Priority Mechanism, alongside an analysis of efficiency that leads to our Dynamic Modular Priority Mechanism. Section 7 concludes. Omitted proofs are relegated to Appendix A.

### v2

@@ Abstract
Priority-based allocation often requires eliminating justified envy, making serial dictatorship (SD) the only non-wasteful direct mechanism with that property. However, SD’s outcomes can conflict with the policymaker’s objectives. We introduce visible fairness, a framework where fairness is evaluated using coarser information. This is achieved by designing message spaces that strategically conceal information which could render desired allocations unfair. We characterize these mechanisms as generalizations of SD, establish conditions for strategy-proofness, and show how to implement distributional constraints. This creates a new trade-off: achieving distributional goals may require limiting preference elicitation, forgoing efficiency gains even when compatible with the constraints.
@@ 1 Introduction
Priority-based assignments are pervasive in a wide range of real-world matching contexts, including university admissions, public-sector placements, and cadet-branch allocations in military academies (see, e.g., Balinski and Sönmez, 1999; Sönmez and Switzer, 2013). In a typical priority-based system, participants are strictly ranked—based on exam scores or a merit list for instance—and are assigned to positions accordingly. The central fairness requirement in these contexts is that no lower-ranked participant should occupy a seat that a higher-ranked participant strictly prefers; otherwise, the latter has a legitimate grievance, known as justified envy (Abdulkadiroğlu and Sönmez, 2003). Such fairness concerns become even more pronounced when the priorities at stake represent strongly protected interests—like property rights or national exam rankings—where even a single instance of justified envy can trigger legal and administrative challenges. [^1] One example of these legal challenges is the Federal University of Bahia hiring suit (Brazil, 2025), where a federal judge blocked the university from hiring a lower-scoring quota applicant and ordered the single vacancy awarded to the exam’s top scorer (Agência Estado, 2025). Another was the Italian national residency “fiasco”, in which the Regional Administrative Tribunal of Lazio annulled a ruling that forced higher-ranked doctors to forfeit more-preferred specialties while lower-ranked peers advanced (Focus.it, 2016). Both rulings treat the harm as a breach of the merit order—i.e., a violation of justified envy—showing that such breaches readily provoke litigation.
Under the standard approach of designing direct mechanisms—where each participant reports a complete ranking over positions—Serial Dictatorship (SD) is in fact the only mechanism that can satisfy non-wastefulness—i.e., not leaving desirable positions unfilled—and no-justified-envy. In SD, the highest-priority participant picks their top choice, the next participant picks from the remaining positions, and so on. This procedure prevents any lower-ranked participant from ending up in a spot that a higher-ranked participant strictly prefers, ensuring no-justified-envy. However, SD can produce allocations that are misaligned with policy goals, such as excessive clustering of top-ranked participants in a small set of elite locations, or undesirable regional or demographic distributions. When considering standard direct mechanisms, there is no alternative to SD that both respects strict priorities and operates on full preference lists. Thus, policymakers appear to face a dilemma: given the requirement for “fairness by priority”, how can one construct rules in pursuit of better distributional outcomes?
@@
1. In the Indian Administrative Services (IAS), officers were formerly assigned to state cadres in a priority-driven process aligned with exam-based merit. This arrangement, which was essentially a serial dictatorship with some modifications, produced undesirable allocations exhibiting homophily, i.e., a propensity for officers to serve in or near their home regions. Such geographic clustering was seen as compromising the national integration objective of the service (Thakur, 2023). A 2017 reform imposed a zone-based scheme: all cadres were partitioned into five geographic zones, and each officer now submits a separate ranking of cadres within each zone rather than a single list over the entire country (see Figure1). The revised mechanism guarantees that no lower-ranked officer receives a cadre preferred by a higher-ranked officer inside the zone where they are ultimately matched, while the zonal structure itself allows for officers to be more evenly dispersed across India—advancing distributional goals without overriding the preferences participants actually report.
@@
>>Zone 2 Uttar Pradesh
@@
>>Zone 5 Andhra Pradesh
@@
2.  Under the U.S. Military Academy matching process of cadets to military branches used in 2006, each cadet (i) ranked the branches and (ii) stated, for every branch, whether they would accept a longer service obligation in exchange for a priority boost (Sönmez and Switzer, 2013; Greenberg et al., 2024). This elicitation did not allow them to report cross-branch trade-offs. Omitting this information let the Academy honour the official order-of-merit list while making it possible to steer more cadets into longer commitments that would have been rejected under full preference elicitation. As Sönmez (2024) explains, this design kept such priority violations hidden:
@@
As the preceding cases show, limiting what participants may report can hide genuine priority breaches. Following this, we say a mechanism is visibly fair when, given the elicited (partial) preferences, no outcome appears to violate priority. For example, if seats are partitioned into zones and participants may only rank seats within each zone, allocating by priority inside every zone looks perfectly fair—even though a cross-zone comparison (never elicited) might reveal a lower-ranked participant holding a seat a higher-ranked participant prefers. By restricting the scope of reported preferences, policy makers can pursue goals such as geographic diversity while keeping any latent violations invisible. A closely related idea appears already in Greenberg et al. (2024), who introduced the notion of detectable priority reversals, a concept that corresponds precisely to visible (un)fairness in the context of the US Army’s branching mechanism.
Inspired by these observations, we examine the design problem of assigning officers to positions under a strict priority ordering while maintaining visible fairness. In contrast to standard models that fix a preference-reporting format, in this framework the policy maker chooses both the message space (the form of partial preferences agents can report) and the outcome rule. Our analysis provides a framework and results on how to configure these two elements together so as to achieve desired policy objectives.
@@
Our analysis delivers three main sets of results. First, we pin down the precise structure that visible fairness imposes on allocation rules. Theorem 1 shows that any visibly–fair mechanism must operate as an $m$-queue allocation: officers are processed in strict priority order, and each is assigned a state that is undominated, among the remaining states, within the partial ranking she is allowed to report. When the message space induces a partition of the state space into zones, this characterization yields the more specific results in Theorems 2 and 3. In this setting, the only visibly-fair rules are partitioned priority mechanisms. Moreover, when rankings are permitted across “zones,” the only visibly-fair rules are ranked-partitioned priority mechanisms. While visible fairness implies serial dictatorship (and therefore strategy-proofness) when using direct mechanisms (Corollary 2), that is not the case for general message spaces. We show in Theorem 4 that strategy-proofness is obtained exactly when the mechanism also satisfies expressiveness and (weak) availability, two properties that rule out profitable deviations when the mapping from message profiles to outcomes is more general.
Second, we introduce a flexible way to encode distributional goals through modular upper-bounds. A quota system is modular when every bound caps groups of officers within an arbitrary subset of states (Definition 10). Modular bounds induce zonal message spaces with a partition of states—formally captured in Definition 13—where all states subject to the same collection of caps fall in the same “zone.” Building on this structure, the Modular Priority Mechanism (Definition 14) processes officers by priority while dynamically clogging zones whose relevant caps have just filled. Theorem 6 proves that this mechanism simultaneously respects every modular bound, remains visibly fair, and is strategy-proof.
Yet, as Example 6.1 and the impossibility result in Theorem 7 shows, in general, no static (one-shot) rule can simultaneously (i) satisfy visible fairness, (ii) respect modular upper-bounds, and (iii) not be Pareto-improved by allocations that respect the caps. The root of the conflict is informational. To guarantee that every admissible report keeps the quotas intact, the policy maker must prune the message space in advance, excluding comparisons whose truthful revelation could otherwise force a violation. This secures fairness and feasibility but withholds preference information that would uncover (and implement) mutually beneficial swaps still compatible with the same caps.
@@
A foundational theory for mechanisms with restricted message spaces comes from Green and Laffont (1986), who study settings where participants are restricted to a limited message space that depends on their true state. Their key insight is that by constraining the information a participant can reveal, the set of implementable outcomes can be expanded beyond what is possible in standard direct mechanisms.
@@
Beyond the already discussed cases of the Indian Administrative Service (IAS) cadre allocation (Thakur, 2023), the U.S. Military Academy’s cadet-branch matching (Sönmez and Switzer, 2013; Sönmez, 2013; Greenberg et al., 2024), and Chinese college admissions (Hu et al., 2025), many real-world mechanisms also limit the extent of preference reporting. In school choice, for example, some systems cap the number of schools an applicant may rank (Haeringer and Klijn, 2009a; Calsamiglia et al., 2010), while others allow applicants to “bundle” schools into groups without inter-group comparisons (Huang and Zhang, 2025).
@@
Dynamic procedures often also restrict preferences and by doing so, might reduce complexity (Pycia and Troyan, 2023). Bó and Hakimov (2022) propose Iterative Deferred Acceptance by letting participants choose from menus, obtaining stable results without demanding complete rankings, and Bó and Hakimov (2024); Mackenzie and Zhou (2022) extend this idea to mechanisms that sequentially offer feasible outcomes—enhancing privacy and performing well in controlled experiments. Even small constraints, like limiting the length of rank-ordered lists, can disrupt classical incentive properties: Haeringer and Klijn (2009b) show how capping the number of ranked schools compromises the usual strategy-proofness of the Deferred Acceptance procedure. Meanwhile, Caspari and Khanna (2025) propose precise conditions for stability and incentives with non-standard preference formats. Collectively, these studies highlight how restricting preference elicitation can open new design possibilities while preserving key desiderata — an insight we leverage in defining and deploying visible fairness. Decerf et al. (2024) is perhaps the most conceptually related work to ours. Their notion of incontestable assignments describes an environment where participants cannot fully observe others’ preferences or placements, leaving them unable to identify certain envy issues. This parallels our concept of visible fairness, in which certain violations become undetectable. The key difference, however, is that Decerf et al. (2024) derive their informational constraints from the participants’ limited ability to view the full outcome, whereas in our framework, these constraints are intentionally designed by the policy maker. Specifically, in our setup the policy maker restricts what participants can report so as to preclude distributional tensions that might otherwise manifest as visible grievances. [^3] In many real-life aplications, such as the IAS hiring and public sector hiring contests in Italy and Brazil, transparency requirements imply the public disclosure of information such as exam papers, scoring sheets, and even interview recordings. These are not considered private and must be accessible to ensure administrative and social oversight (Autorità Nazionale Anticorruzione, 2025; Controladoria-Geral da União, 2023; Abizada and Bó, 2021). Under these informational circumstances, incontestability might become equivalent to standard elimination of justified envy, and therefore under single priority imply serial dictatorship. In addition, while Decerf et al. (2024) accommodate a variety of school-specific priorities, our model considers a single, strict priority ranking that orders all participants.
Another paper that considers mechanisms using alternative message spaces is Cavallo and Dogan (2024). The authors analyze Italy’s nationwide teacher‑mobility scheme, in which teachers may rank entire municipalities, districts, or provinces—nested geographic units that bundle many schools into a single item on their list. They show that the tie‑breaking rule used to resolve these coarse rankings might allocate lower‑priority teachers ahead of higher‑priority ones, creating detectable instances of justified envy. They also show that these result in legal challenges: Italian courts have repeatedly upheld merit‑based claims, and parliamentary testimony records more than 1,000 lawsuits filed each year, on average, over such priority violations.
@@
Partial preferences have also been studied in other contexts within market design. One strand lets participants declare indifference classes directly: Erdil and Ergin (2017); Manjunath and Westkamp (2021); Andersson et al. (2021) build mechanisms that treat weak orders — strict ranks punctuated by ties — as the primitives, and then exploit those ties to recover efficiency and strategy‐proofness. A second strand considers problems in which not every pair of outcomes can be compared, evaluating which adaptation of standard properties, such as stability, can nonetheless sustain strategy-proofness under suitable conditions (Caspari and Khanna, 2025; Kuvalekar, 2023). Typically, these frameworks rely on “weak stability,” where participants who are indifferent or indecisive simply cannot block assignments. In contrast, our approach presumes participants do have complete preferences but are deliberately constrained from revealing them in full.
@@
At the core of this field lies the tension between rigid distributional quotas and the flexible preferences of participants. Fragiadakis and Troyan (2017) demonstrated this through military cadet matching, where static reservation systems created inefficiencies by locking seats for specific groups prematurely. Their dynamic quota mechanism represented a paradigm shift, adjusting reservation targets based on revealed preferences while maintaining strategy-proofness. This approach inspired subsequent innovations like the Adaptive Deferred Acceptance (ADA) mechanism by Goto et al. (2017), which introduced hereditary constraints—rules where satisfying a constraint automatically satisfies all its subsets. The ADA mechanism’s success in Japanese medical residency matching showed that carefully designed constraints need not sacrifice core market principles like nonwastefulness and strategy-proofness.
Structure of the paper. In Section 2, we introduce the model and definitions, covering partial preferences, feasible allocations, and our message-space framework. Section 3 then characterizes visibly fair mechanisms, identifying them as queue-allocation variants and examining applications such as zonal message spaces. Section 4 turns to incentives, specifying exactly when these mechanisms are strategy-proof via the conditions of expressiveness and availability. Sections 5 and 6 present our results on distributional objectives, including modular upper-bounds and the Modular Priority Mechanism, alongside an analysis of efficiency that leads to our Dynamic Modular Priority Mechanism. Section 7 concludes. All proofs are relegated to Appendix A.